### v3

# Micellization in active matter of asymmetric self-propelled particles: Experiments

Abstract
Active matter composed of self-propelled particles features fascinating self-organization phenomena, spanning from motility-induced phase separation to phototaxis to topological excitations depending on the nature and parameters of the system. In the present paper, we consider micelle formation by active particles with a broken symmetry having a circular back and a sharpened nose toward which the particles accelerate. As we demonstrate in experiments with robotic swarms, such particles can either remain in the isotropic phase or form micelles depending on the location of their center of inertia, in accordance with a recent theoretical proposal [T. Kruglov and A. Borisov, Presentations and Videos to 7th Edition of the International Conference on Particle-based Methods (2021), Vol. CT07, p. 2]. Such a behavior is observed for both nonchiral particles moving linearly and placed in a parabolic potential and for chiral particles moving along circular trajectories on a flat surface. By performing experiments with single robots and two-robot collisions, we unveil that the observed emergence of micellization associated with shifting robots’ center of inertia towards their noses is governed by at least two-particle effects, in particular, by a difference in the formation of stable two-robot clusters. Finally, we consider the dependence of micelle lifetime and formation probability as well as two-robot collisions on friction between the lateral surfaces of the robots. Crucially, the predicted micellization does not involve any solvation shells that give rise to the micellization of surfactants but is instead driven by an interplay of activity and particle shape asymmetry.
I Introduction
Large assemblies of particles that can self-propel or self-rotate by converting internal or ambient energy resources to directed motion demonstrate the emergence of collective phenomena Reynolds (1987); Vicsek et al. (1995) and are referred to as active matter Vicsek and Zafeiris (2012). Such systems span the entire range of soft condensed matter, from tissues Balasubramaniam et al. (2021) and bacterial colonies Pierce et al. (2018); Be’er et al. (2020) to colloidal particles Bricard et al. (2013) or even collections of simple moving robots Giomi et al. (2013); Barois et al. (2019); Patterson et al. (2020); Li et al. (2021); Boudet et al. (2022). There is a rich variety of self-organization phenomena and clustering effects in active matter systems, including motility-induced phase separation Cates and Tailleur (2015), the formation of colloidal crystals Palacci et al. (2013); Ginot et al. (2018); Mousavi et al. (2019), the emergence of chiral edge states Yang et al. (2020); Shankar et al. (2022); Souslov et al. (2017), and diverse physics of topological defects Duclos et al. (2020), to name a few.
However, the formation of micelles—round or spherical-shaped assemblies of elongated particles in which they orient one of their nonequivalent edges to the inner region of a cluster and the other one to the outer region—has been demonstrated only in Janus particles covered with surfactants Zhang et al. (2015). This phenomenon appeared elusive in self-propelled particles without charge displacement, until the recent theoretical proposal of active - matter micellization driven purely by particle shape asymmetry Kruglov and Borisov (2021) and experimental studies of rotelles, the rotational analogs of micelles formed by surfactant-like chain of self-rotating active particles Scholz et al. (2021).
In the present paper, we consider ensembles of self-propelled particles that have the teardrop shape shown in Fig. 1(a). Particles having a similar shape occur in biological systems, for example, specific types of blood cells Gütgemann et al. (2014), and can be fabricated on the millimeter to micrometer scale with the help of various techniques, e.g., self-propelled asymmetric Janus particles in Refs. Dai et al. (2016); Shah et al. (2021); Liu et al. (2022). To demonstrate experimentally that such particles can form micellelike clusters depending on the center of inertia location at the particle axis, we construct a swarm of teardrop-shaped vibrating robots (bristle-bots) shown in Figs. 1(b) and 1(c) which are based on the Swarmodroid 1.0 platform Dmitriev et al. (2023). Such robots move with a controlled level of activity and are placed in a shallow parabolic potential of a satellite dish to prevent condensation at the system boundary, typical of self-propelled particles Giomi et al. (2013); Deblais et al. (2018); Boudet et al. (2021), Fig. 1(a), or on a flat surface with their trajectories modified to be slightly chiral (corresponding to robots moving along circular loops). In accordance with the predictions of Ref. Kruglov and Borisov (2021), we report micelle formation both for swarms on a flat surface and in a parabolic potential and address the dependence of their formation probability on the packing density of robots, their center of inertia location, velocity of their motion, and the friction coefficient between lateral surfaces of robots. Finally, we uncover the microscopic mechanisms responsible for such a sharp dependence of the micellization on the location of individual robots’ center of inertia by performing experiments with single robots as well as with two-robot collisions in different conditions.
The paper is organized as follows. In Sec. II, we describe our experimental platform. Then, in Sec. III a key quantitative parameter is introduced which we use to detect micellization. Section IV presents the results of experiments with robotic swarms and includes three subsections. Section IV.1 includes studies of the condensation of robots at the boundary when they propel along straight trajectories on a flat surface. The next Sec. IV.2 features a demonstration of robots’ micellization on a flat surface for robots moving along chiral trajectories and includes the study of the micellization statistical properties. Then, Sec. IV.3 considers micellization of robots in a parabolic potential, including the dependence of the lifetime and formation probability of micelles on the packing density of robots, their motion velocity, and the friction between their lateral surfaces for different locations of the robots’ center of inertia. The experiments considering microscopic origin of the observed micellization transition are discussed in Sec. V. Section V.1 includes studies of individual robot motion in a parabolic potential for different center of inertia locations demonstrating that the behavior at a single-robot level does not demonstrate any pronounced changes. Section V.2 features the studies of two-robot collisions pointing toward a considerable increase in the formation of stable quarter-micelle clusters, hinting towards its possible influence on the emergence of the micellization. Section VI contains the final remarks and discussion of the obtained results.
II Experimental platform for active matter: asymmetric robots
In our experiments, we implement the teardrop-shaped particles as self-propelled bristle-bots converting their vibration to a directed motion, Figs. 1(b)- 1(e). Such robots consist of a printed circuit board [Fig. 1(d)] carrying a vibration motor, a battery, and circuitry for infrared remote control, which allows turning the robots on and off simultaneously and varying their vibration activity (i.e., a self-propulsion velocity) by changing the pulse width modulation (PWM) duty cycle of the motor voltage, and a three-dimensional (3D) printed body with elastic bristles at the bottom [Fig. 1(e)] Dmitriev et al. (2023). The system dynamics is extracted using a tracking pipeline based on our custom recognition software Dmitriev et al. (2023) and the use of ArUco markers.
The length of each robot is , the diameter of the round part is , and the nose angle is corresponding to particles in a complete micelle. We choose such a value of to clearly distinguish micellization from crystallization, as the point symmetry is incompatible with a crystalline order. The height of the robots including the bristles is mm. The robot’s center of inertia lies on its mirror symmetry axis at the distance from its nose cusp. With the help of an additional load, we are able to move the center of inertia from mm (point ), which is the center of inertia of a robot without additional load [Fig. 1(b)], to mm (point ) for a robot with a load [Fig. 1(c)]. We attach a load consisting of an M616 screw and three M6 nuts made of steel, which are fastened to a hole near the robot’s nose, to move the center of inertia to point . In the following, we denote such robots as -robots. To increase the mass of a robot by the same amount while maintaining the center of inertia at point , we instead add a load consisting of a steel DIN125A M18 washer, which is glued to the robot cap at point Sup . We denote these robots as -robots. The mass of -robots and -robots is the same and equal to g. We study swarms composed of , , , and robots.
Figures 1(f) and 1(g) show the velocities of - and -robots, respectively, as dependencies of the PWM duty cycle, which represents the vibration activity. Due to the complex dynamics involved in the conversion of vibration to directional motion by bristles Giomi et al. (2013); Scholz et al. (2016); Porvatov et al. (2021); Dmitriev et al. (2023), there is a distribution of velocities between different robots at each value of the vibration activity, which we show as a color-encoded histogram. The standard deviation of this distribution reaches one third of the mean value for -robots at the vibration activity of . Therefore, we do not refer directly to the robot velocity throughout the paper, but instead use vibration activity which is exactly reproducible. It is seen that, despite their different weight distributions, the velocities of -robots and -robots coincide for vibration activities up to that are studied in the following.
III Micellization parameter
To quantify the formation of micelles, we introduce the micellization parameter
where and are the in-plane coordinates of the noses of particles with indices and , is the number of particles in a complete micelle (in our case, ), is the characteristic distance between the noses of particles at which they are considered forming a micelle (here, we choose mm), and is the total number of particles in the system. Figure 2(a) shows the values of such micellization parameter extracted experimentally for several model systems: an isolated complete micelle with , several types of isolated incomplete micelles with , an inverse micelle with , a structure of densely packed robots with crystalline order, a single robot, and two types of two-robot clusters (“tandem” and “heart”) that, as will be shown in Sec. V, are important for understanding the processes facilitating micelle formation. The terms in the sum by index tend to unity for a complete micelle while rendering values orders of magnitude smaller for other configurations. However, due to the exponential dependence on the distance between the robots’ noses, the micellization parameter yields values somewhat lower than unity for experimentally observed micelles, with the majority of observed micelles corresponding to . For incomplete micelles and nonmicellar structures, the micellization parameter yields exponentially low values. The normalization prefactor ensures the invariance of the parameter with respect to the number of robots. For fully chaotic systems, the micellization parameter yields the value independently of the number of robots in the system. The same value corresponds to the system consisting of a single robot and is the lowest possible value of the micellization parameter.
Figure 2(b) shows the numerically calculated dependencies of the micellization parameter on the number of particles in the system up to . The fully micellized systems [the solid red lines in Fig. 2(b)] correspond to the particles that are added to the system sequentially to form complete micelles. In this case, the micellization parameter is at its maximum for the numbers of particles that are multiples of eight, with intermediate values corresponding to the presence of incomplete micelles resulting in the lower values of . The curves for a single micelle among a disordered environment [the black solid lines in Fig. 2(b)] are obtained by adding randomly oriented particles at random coordinates within a box around a complete micelle. The coordinates and angular orientations of the added particles are distributed uniformly. In this case, the micellization parameter decreases as with the number of particles. Therefore, the parameter introduced in Eq. (1) characterizes the fraction of micelles in the swarm.
We also study the dependence of the micellization parameter on the distance between the particles in a single micelle by introducing random shifts of the particles’ noses from the center of a micelle, Fig. 2(b). The components of the shift vectors are normally distributed with the zero mean and the dispersion that varies from to mm (approximately of the robot’s diameter). The micellization parameter decreases exponentially depending on this amplitude, producing the same value for a complete micelle with mm gaps between the robots and an ideal incomplete micelle without one robot.
IV Experimental studies of micellization in robotic swarms
IV.1 Boundary condensation of non-chiral robots on a flat surface surrounded by a vertical wall
| -robots | ||||||||||||
| № | 1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 | 9 | 10 | Avr, s | Msd, s |
| t, s | 83 | 76 | 83 | 69 | 21 | 17 | 10 | 52 | 24 | 70 | 49.6 | 24.6 |
| № | 11 | 12 | 13 | 14 | 15 | 16 | 17 | 18 | 19 | 20 | ||
| t, s | 74 | 48 | 29 | 47 | 46 | 14 | 24 | 69 | 79 | 57 | ||
| -robots | ||||||||||||
| № | 1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 | 9 | 10 | Avr, s | Msd, s |
| t, s | 22 | 29 | 32 | 35 | 47 | 48 | 12 | 27 | 41 | 27 | 33.3 | 12.9 |
| № | 11 | 12 | 13 | 14 | 15 | 16 | 17 | 18 | 19 | 20 | ||
| t, s | 37 | 18 | 30 | 26 | 32 | 25 | 63 | 26 | 62 | 27 | ||
We start with considering a swarm of teardrop-shaped robots moving on a flat surface and confined by a circular barrier (Table 1). The barrier is 3D printed with PLA plastic, and its diameter is cm; see Fig. 3. As seen in Figs. 3(a)-3(d) and Fig. 3(e)-3(h), the robots demonstrate a pronounced condensation at the barrier for both considered center of inertia locations, with characteristic condensation times s for -robots and s for -robots, respectively, see supplementary video 1. These values are obtained by averaging over experiments for each type of the robots starting from a randomly arranged swarm. The detailed results are provided in Table 1. In experiments, the measured value of condensation time corresponds to the time frame when all robots localize at the barrier for the first time, with the only allowed exception of one or two robots roaming along closed circular trajectories in the bulk due to the chirality that is inevitably present in their motion.
Once condensed at the boundary, the robots cannot return to the center, which hinders the further emergence of large clusters in the bulk. Moreover, the robots located at the barrier touch it with their noses or lateral surfaces; hence, an axially symmetric micelle cannot form at the barrier. Similar condensation has already been reported Giomi et al. (2013) and even applied to engineer specific swarm behavior Deblais et al. (2018) or collective functionalities Boudet et al. (2021). Its appearance appears absolutely intuitive and is related to the negative curvature of the boundary. As shown in Secs. IV.2 and IV.3, the characteristic times of a single micelle formation are much longer than the condensation times observed in Fig. 3.
As a result, in order to study micellization, such boundary-assisted condensation of robots should be avoided for as long as is feasible, which can be achieved with several strategies.
1.
Considering abrupt barriers of the same type, but with diameters large enough for the system to micellize prior to the condensation is achieved. Indeed, the characteristic time of the bulk micellization transition should be independent of the systems’ boundary, while the condensation time grows with the barrier diameter (considering fixed filling density, since it is defined by the mean distance each particle should travel to collide with the barrier).
2.
Repulsive boundaries, for example, creating air flows in a similar manner to those described in Ref. Li et al. (2021) or incorporating mechanically moving parts to orient the robots’ noses back to the center of the experimental area.
3.
Passive barriers with specific shapes that facilitate returning robots to the bulk after traveling a certain distance along the boundary. For example, Refs. Deseigne et al. (2010); Kumar et al. (2014) consider sunflower-shaped barriers used together with systems of self-propelled disks or millimeter-sized tapered rods, respectively.
4.
Chiral trajectories of robots. If each robot moves along closed circular trajectories with diameter much larger than the characteristic robot size , yet much lower than the barrier diameter , then the robots can be considered (locally) self-propelling forward, yet they can localize in the bulk and touch the barrier only as a result of collisions with other robots. Moreover, the robot can still return to the bulk after condensing at the barrier if its motion direction along the barrier coincides with its own chirality. Such chiral systems are considered in Refs. Deblais et al. (2018); Barois et al. (2020); Boudet et al. (2021); Siebers et al. (2023).
5.
Soft confining potentials preventing condensation. Robotic swarms in external potentials have been considered in the case of an inclined plane with phototactic behavior of robots whose activity is defined by the level of external illumination Ben Zion et al. (2023) and even in more complex cases of robots that can sense and alter the illumination pattern on the surface they travel on Wang et al. (2021).
| № | 1 | 2 | 3 | 4 | 5 | 6 | 7 | 8 | 9 | 10 |
| 173–174 s | 7–8 s | 174–176 s | 54–55 s | |||||||
| № | 11 | 12 | 13 | 14 | 15 | 16 | 17 | 18 | 19 | 20 |
| 22–23 s | 48–49 s | 80–83 s | ||||||||
| 251–252 s | ||||||||||
In the following, we consider two strategies: chiral robots moving along circular trajectories in clockwise or counterclockwise directions that allow us to demonstrate the emergence of micellization in robotic swarms on flat surfaces, and non-chiral robots in a parabolic potential.
IV.2 Micelle-like clusters in swarms of chiral robots on a flat surface
We start with considering swarms with an even number of robots in which half of the robots have clockwise (CW) chirality, i.e., move along the circular trajectories centered to the right from the robot’s propulsion vector direction, while the rest half robots move in the opposite, counterclockwise (CCW) direction, Fig. 4(a). The robots are made chiral by adding an additional inclination of the bristles in the planes of their arrays. We select the opposite additional inclination angles for the front and bottom bristle arrays, thus adding the rotational component to the robots’ propulsion. In experiments, we set the number of robots rotating clockwise and counterclockwise equal to zero out the net chirality of the system and avoid chirality-assisted effects such as the formation of unidirectional boundary flows Yang et al. (2020); Barois et al. (2020). As outlined in Sec. IV.1, the robots’ trajectories are tuned in such a way that the radius of the trajectory along which the robot moves is less than the radius of the barrier, but is larger than the size of the robot, which allows us to consider the robot’s motion locally linear. Moreover, the chirality of the robot’s trajectory allows it to turn away from the barrier in certain situations, and thus additionally prevents the condensation mentioned in Sec. IV.1. The considered chiral robots are placed on a flat surface made of plywood and are bounded by a plastic barrier with diameter cm.
In our experiments, we vary the location of the center of inertia considering two cases: -robots [Figs. 4(b)-reffig:Chiral(e)] and -robots [Figs. 4(f)-4(i)]. We capture the collective dynamics of the robotic swarm on video for five minutes and repeat the experiment times for each robot type, considering different randomly chosen initial locations of the robots. In all experiments, we set the high robot activity ().
As seen in Figs. 4(b)-4(e), for -robots micelle formation is not observed, although in Fig. 4(c) it is clearly seen that the robots turn away from the boundaries of the barrier without condensing on it (see the red and blue tracks highlighting the trajectories of CW- and CCW-chiral robots, respectively). In contrast, for -robots, a micelle is formed at s, Fig. 4(h). However, it decomposes in several seconds after the formation and is not present in Fig. 4(i), suggesting general instability of micelles in the system with chiral robots. Both cases are illustrated in supplementary video 2.
The reason behind such an instability is the following. In a two-robot cluster (representing a quarter of a micelle) composed of a CW-chiral robot placed to the left from a CCW-chiral robot (looking towards the direction of its propulsion vector), two such robots possess propulsion components that attract them together while they propel forward in the same direction. As a result, such a cluster retains its shape during motion. The formation of this type of two-robot cluster is analyzed in more detail in Sec. V.2. However, if we instead put the CW-robot to the right of the CCW-robot, then repulsive propulsion components appear that pull the robots apart. Therefore, such a cluster cannot exist for a long time. The stability of analogous configuration with two CW- or two CCW-chiral robots will, obviously, be intermediate between the two considered cases. The same logic applies to configurations with larger numbers of robots, i.e., half-micelles, complete micelles, and other clusters. In addition, every cluster can be composed with several different combinations of CW- and CCW-chiral robots, for example, the micelle in Fig. 4(h) is composed from three CW- and five CCW-chiral robots. The number of these combinations increases with the number of robots in the cluster, as is known from combinatorics, and only part of these configurations are stable.
Next, we analyze the statistical properties of the micelles observed in the discussed experiments with swarms of chiral robots on a flat surface. As seen in Table 2, no micelles were observed in the system with -robots. However, in the case of -robots, a micelle is formed in of the experiments, with the mean lifetime s limited by the instability of the micelles formed by chiral robots. The insets of Table 2 highlight CW-chiral robots with red and CCW-chiral robots with blue, showing different ratios of CW- and CCW-chiral robots in the observed micelles. Finally, in accordance with our initial considerations, the characteristic time of micelle formation s significantly exceeds the typical condensation time s (for -robots, see Sec. IV.1) significantly, justifying the necessity of considering chiral trajectories.
To summarize, condensation at the vertical wall barrier is significantly reduced in swarms of chiral robots moving on a flat surface compared to nonchiral robots. Moreover, micelle formation is observed only in the case of -robots. However, a quantitative analysis of the micellization is complicated by the short micelle lifetime of just a few seconds, making it necessary to consider various strategies toward its increase.
IV.3 Micellelike clusters in swarms of nonchiral robots in a parabolic potential
We proceed with studying a system of nonchiral -robots moving with low activity ()[Fig. 5(a)] and placed in a parabolic satellite dish having the dimensions of [Fig. 1(a)]. In contrast to a hard wall, such a dish creates a soft localizing potential and, as we will demonstrate further, prevents the condensation of robots at the boundary. The micellization parameter fluctuates during the system evolution but does not demonstrate any pronounced growth characteristic of micellization, Fig. 5(a). In this case, robots are predominantly directed away from the vertex of the parabolic potential Dauchot and Démery (2019). Covering the lateral surfaces of robots with abrasive tape in order to increase the friction between them does not result in any systematic changes of the micellization parameter, see the blue solid line in Fig. 5(a).
The situation changes qualitatively for -robots, Fig. 5(b). Initially randomized, the system first demonstrates a low micellization parameter corresponding to the absence of micelles (see the inset for showing the robots’ distribution), followed by an increase in the micellization parameter corresponding to formation of incomplete micelles (see the inset for ). Finally, at , a plateau of the micellization parameter is observed, corresponding to the formation of a stable micelle shown in the inset at , as demonstrated in Supplementary Video 3.
Next, we consider the dependencies of the micellization parameter on the activity of the robots, their center of inertia location, and the packing density that is varied by changing the number of robots in the parabolic potential. For systems of -robots with low () and high () activity shown in Figs. 6(a) and 6(b) the micellization parameter fluctuates near the minimal level () independently of the density of the system, and micelle formation is not observed; see Table 3. -robots do not demonstrate any micelles as well in the experiments with sparse swarms of robots; see Figs. 6(c) and 6(d). However, stable half-micelles appear in these systems, characterized by an increase in the micellization parameter (). The formation of transient micelles with a short lifetime of a few seconds emerges at the higher density , and is well expressed for , indicated by multiple peaks of the micellization parameter in Figs. 6(c) and 6(d) both for low () and high () activities. Increasing robot activity results in shorter average micelle lifetimes, namely s for and s for for high robot activity, compared to s for and s for for low activity.
Next, we modify the system by covering the lateral surfaces of all robots with an abrasive (P800 sandpaper) Sup to consider how the micellization changes in the case of high friction between the particles, which may prove important for microscale implementations. As seen in Figs. 6(e) and 6(f), the behavior of -robots is similar to the case of particles with low lateral friction; Figs. 6(a) and 6(b). The micellization parameter slightly fluctuates, indicating the absence of micellization. The same is confirmed by the data in Table 3.
| Smooth lateral surfaces | ||||||||||||
| Low activity | High activity | |||||||||||
| 1 | 69–77 s | 224–226 s | 70–71 s | 71–75 s | ||||||||
| 121–123 s | 102–106 s | 116–121 s | ||||||||||
| 127–131 s | ||||||||||||
| 229–230 s | ||||||||||||
| 248–251 s | ||||||||||||
| 284–285 s | ||||||||||||
| 2 | 142–146 s | 106–116 s | 59–60 s | 66–75 s | ||||||||
| 186–198 s | 147–149 s | 222–225 s | 102–105 s | |||||||||
| 212–214 s | ||||||||||||
| 3 | 267–268 s | 184–185 s | ||||||||||
| 295–300 s | ||||||||||||
| 4 | 192–204 s | 157–166 s | 41–43 s | 69–70 s | ||||||||
| 189–243 s | 192–193 s | 193–194 s | ||||||||||
| 5 | 53–54 s | 41–48 s | ||||||||||
| 99–101 s | ||||||||||||
| Abrasive lateral surfaces | ||||||||||||
| Low activity | High activity | |||||||||||
| 1 | 169–176 s | 34–47 s | ||||||||||
| 265–273 s | ||||||||||||
| 287–300 s | ||||||||||||
| 2 | 216–232 s | 229–234 s | 41–43 s | |||||||||
| 248–252 s | ||||||||||||
| 3 | 217–224 s | 40–44 s | ||||||||||
| 4 | 197–233 s | 62–300 s | 93–94 s | |||||||||
| 5 | 222–224 s | 234–241 s | ||||||||||
| 263–270 s | ||||||||||||
However, for -robots with abrasive lateral surfaces, the situation becomes strikingly different compared to the case with smooth lateral surfaces. As seen in Table 3, micelle formation for low activity () is observed only for the system with , while for sparse swarms with or, in contrast, dense swarms with the micelles are not observed. Moreover, for high activity (), micelles are observed most frequently for intermediate density , while for micelle formation is less pronounced and appears almost completely suppressed for , with only a single experiment featuring the formation of a complete micelle. Instead, distorted micelle-like clusters consisting of two half-micelles shifted a few centimeters relative to each other frequently form in these systems, characterized by plateaus of the micellization parameter with values ; see Figs. 6(g) and 6(h). The average lifetimes of micelles are s for (excluding the anomalously stable micelle discussed separately in Sec. VI), s for , and s for , significantly exceeding those for the micelles formed by smooth robots. Decreasing robot activity leads to an even longer average micelle lifetime of s for . Thus, it is seen that, in contrast to robots with smooth lateral surfaces, micelles are more stable in swarms of abrasive robots, and increasing robot activity significantly enhances the probability of micelle formation.
This difference in micellization between smooth and abrasive robots could be related to the improved stability of the clusters formed by robots with high lateral friction. Systems of robots with smooth lateral surfaces demonstrate fewer micelles at lower densities, while at higher densities there are more micelles that decompose quickly. This might be explained by the increased probability of micelle formation during the decomposition of other clusters, which are generally characterized by low stability. With high lateral friction, micelles demonstrate longer lifetimes, apparently related to restriction of the relative lateral movement of adjacent robots in a cluster. This would presumably lead to increased micellization in systems forming clusters that are more likely to merge into micelles, which, once formed, would remain stable. In contrast, for systems where disordered clusters are more likely to form, their enhanced stability would lower the chances of decomposition followed by micelle formation, resulting in decreased micellization. Such a variety of possible mechanisms emphasizes the complex interplay between the asymmetry of individual particles, their activity, and friction in the considered system, which may be crucial for microscale realizations, for example, using Janus particles Zhang et al. (2015); Hils et al. (2021).
To evaluate the interaction of micellelike clusters with nonmicellar aggregates, lifetimes of manually assembled micellelike clusters were assessed. First, manually assembled micellelike clusters in the absence of nonmicellized robots were examined. They remained stable for the entire duration of the experiment (several minutes), independently of the vibration activity of the robots (low or high), their lateral friction (smooth or abrasive side surfaces), and even of the center of inertia location-micellelike clusters composed of both -robots and -robots remained stable (see supplementary video 7). From this we can conclude that micellization is governed by the probability of micellelike cluster formation, but once formed, micellelike clusters are stabilized by the parabolic potential. After that, nonmicellized robots were added to evaluate the interaction of micellelike clusters with other aggregates and individual robots. These systems most often displayed two kinds of behavior: micellelike clusters either decomposed in the first few seconds after a collision with surrounding robots, or remained stable for the entire experiment if such collisions did not occur (see supplementary video 8). Additionally, -robots occasionally formed new micellelike clusters, while -robots did not (see the supplementary material Sup for a detailed analysis). Therefore, micellelike clusters continuously form and decompose with probabilities governed by the interactions between individual robots and various clusters.
V Microscopic mechanisms of micelle-like clusters formation
To unveil the physical mechanisms behind the increased micellization in swarms of robots with the center of inertia located closer to the robots’ noses, we perform a series of experiments with single robots to study how their trajectories vary for different center of inertia locations and consider two-robot collisions to address the changes in their scattering associated with the center of inertia location.
V.1 Similar single-particle dynamics of robots with different center of inertia locations
We start by considering the motion of a single self-propelled robot placed in a parabolic potential implemented with the help of a satellite dish, the same as in the swarm experiments discussed in Sec. IV.3. The dish has two different main axes cm and cm, and therefore its horizontal cross sections are ellipses. Such a soft localizing potential allows us to study the robot’s motion for an arbitrary duration, with the robots neither escaping the area captured by the camera nor condensing at the boundary.
In our experiments, we vary the robot activity (defined by PWM), the center of inertia by setting it at points and , and the initial robot placement configuration defined by the robot’s position and angular orientation. PWM takes the values (the low activity) and (the high activity) corresponding to the robot motion velocities of approximately and for -robots, and cm/s and cm/s for -robots, respectively. Four initial robot placement configurations are considered: (i) the “center” configuration corresponding to the robot placed at the vertex of the parabolic potential and its propulsion vector oriented along the semiminor axis, Fig. 7(a); (ii) the “edge” configuration with the robot placed at the distance cm from the vertex at the semiminor axis and the propulsion vector directed towards the vertex point, Fig. 7(b); (iii) the “orbit CCW” configuration with the robot placed at the semi-minor axis at the distance cm from the vertex and the propulsion vector directed along the semimajor axis in such a manner that the robot moves in a CCW-direction, Fig. 7(c); and (iv) the “orbit CW” configuration when the robot is placed at the distance cm from the vertex at the semi-minor axis and the propulsion vector is directed along the semimajor axis resulting in a CW-directed robot’s motion, Fig. 7(d). For each set of parameters, we repeated the experiment with three different robots whose motion has been recorded during s. To study the kinematic characteristics, we introduce the radius vector of robot , defined in polar coordinates and with the origin at the vertex of the parabolic potential and the angle measured from the major semiaxis of the parabolic satellite dish. The radius vector points to the center of the robot’s marker. We also introduce the robot orientation angle between the robot propulsion vector and the radius vector.
Figures 7(e)-7(l) demonstrates the time evolution of the radius vector components and the orientation angle of a single robot in the parabolic potential. The four previously described initial placement configurations of the robot moving with high activity () are studied for an -robot and an -robot with results displayed in Figs. 7(e)-7(h) and Figs. 7(i)-7(l), respectively. It is seen that there are two major types of individual robot motion, in accordance with Ref. Dauchot and Démery (2019): a car drifting-like climbing motion, whose distinctive feature is the propulsion vector directed outwards the vertex of the parabolic potential and the average velocity directed tangentially to height isolines, and the orbiting motion with the propulsion vector oriented tangentially to the elliptical trajectory of the robot.
The first type of motion is characterized by the orientation angle close to ; see the dashed horizontal lines in Figs. 7(e)-7(l). Moreover, the radial coordinate and the polar angle change slowly (up to some fast fluctuations), with most of the curves corresponding to the robot completing a single revolution around the potential’s vertex during the observed timespan [see the blue, green, and orange solid lines in Fig. 7(e) and the green solid line in Figs. 7(g), 7(k), and 7(l)] or even an incomplete rotation [the orange solid lines in Fig. 7(h) and Figs. 7(i)-7(l) and the green solid lines in Figs. 7(i) and 7(j)]. The radial coordinate changes with the same oscillation period as corresponding to the robot moving at the same height along the elliptical trajectory characterized by different values of the semiminor and semimajor axes. Due to the ellipticity of the trajectory, oscillations with the same period can also be observed in the orientation angle , which would not exist if the trajectory was a perfect circle. The described climbing motion occurs for the -robot, Figs. 7(e)-7(h) [the blue solid line in Fig. 7(e), the green solid lines in Figs. 7(e) and 7(g), and the orange solid lines in Figs. 7(e) and 7(h)] as well as for the -robot [the green and orange solid lines in Figs. 7(i)- 7(l)]. The fast fluctuations mentioned previously, which are clearly manifested in the radial coordinate and the orientation angle , are related to the robot’s oscillating up-and-down motion around a steady height level, see supplementary video 4. In addition, the rattling of with frequencies greatly exceeding those of the oscillations of and an amplitude that may in some cases reach [see Fig. 7(e)] results from a noise in the robot’s propulsion vector orientation due to the bristle-bot vibrating motion mechanism. It is seen that such rattling is only present in the orientation angle but absent in the polar angle .
The second type of single robot motion is characterized by fast oscillations of the radial coordinate accompanied by a sawtooth waveform in the polar angle , both having the same period, Fig. 7. This signifies that the robot repeatedly traverses an elliptical trajectory, while the period of these oscillations is inversely proportional to the robot’s tangential velocity. Moreover, oscillations with a more complicated waveform are observed in the orientation angle that stays either in the vicinity of or , indicating that the robot’s propulsion vector direction is well aligned with its direction of motion along the elliptical trajectory. Thus, in contrast to the climbing motion described above, this second type of motion resembles the movement of a car running on an oval track by proper steering. Similarly to the previous motion type, this orbiting motion is observed for both the -robots shown with the blue and orange solid lines in Figs. 7(f) and 7(g) and the blue solid line in Fig. 7(h), and the -robots shown with the blue solid lines in Figs. 7(i)-7(l).
In some cases, a hybrid type of a single robot motion can be observed, characterized by the orientation angle taking the values between and and the robot making between two and three complete revolutions around the vertex per minute; see the solid green lines in Figs. 7(f)-7(h). The presence of such motion profiles with intermediate characteristics demonstrates that the two previously described regimes may be considered as limiting cases of a single type of motion corresponding to the robot moving periodically along an elliptical trajectory with a certain angle between the robot’s velocity vector and its propulsion vector. Greater values of this angle correspond to more pronounced radial oscillations with a longer period.
To evaluate the role of collisions in the single-particle motion of robots, rotational excitation was applied in the form of instantaneous strikes to various points of the robot side surface far from its center of inertia. The strike causes a transient rotation, which damps in s for -robots and s for -robots. After that, the robots return to their normal movement pattern, that is, climbing or orbiting motion. Detailed analysis of these experiments is provided in the supplementary material Sup . Such damping times are significantly smaller than other characteristic times in the system. Therefore, transport of angular momentum can be neglected due to its almost instantaneous damping by bristle-bots, and robot collisions that do not result in cluster formation are unlikely to play a significant role in micellization.
In summary, the motion profiles of -robots and -robots demonstrate similar properties. It is also seen from the curves for the orbiting motion in Fig. 7 that -robots move along smaller ellipses compared to -robots. Therefore, those of the -robots that move slightly closer to the vertex of the parabolic potential than -robots, are presumably more likely to collide with each other attempting to form clusters, some of which may end up being stable. However, as seen in Fig. 6 and Table 3, micelles do not form in -robot swarms. Thus, the increase in -robots collisions seemingly leads to the formation of clusters that are either unstable, or stable but not contributing to (or even preventing) micelle formation. Therefore, the increased micellization of -robots compared to -robots cannot be explained by differences in the movement of a single particle, and collisions of at least two particles should be considered.
V.2 Dominant stable two-particle cluster type defined by particle center of inertia location
To address microscopic micellization mechanisms beyond the single-particle level, we proceed with considering the collisions of two self-propelled robots for different sets of parameters: low () and high () activity of the robots, abrasive and smooth lateral surfaces, and impact angles ; see Fig. 8(a). Despite single-robot experiments were conducted in a shallow parabolic potential of a satellite dish, we consider nonchiral robots moving linearly on a plane when studying two-robot collisions, as controlling the impact angle for robots moving along elliptical or circular trajectories appears extremely challenging.
Figure 8 shows the key results for - and -robots with smooth and abrasive lateral surfaces moving with low and high activities and colliding at impact angles [Figs. 8(a)-8(d)] and [Figs. 8(e)-8(h)] while additional data for low activity, as well as for are considered in the supplementary material Sup . For each set of parameters, we perform three to four experiments with different pairs of robots to ensure that the observed properties of scattering and cluster formation indeed represent the characteristic behavior of the considered system rather than the specific properties of certain robots. In the following, we focus on the evolution of the distance between the noses of the robots and the angle between their propulsion vectors.
It is clearly observed that, regardless of the impact angle and the presence of an abrasive coating on their lateral surfaces, -robots form stable (i.e., moving without any considerable shape variations) heart-shaped two-robot clusters after a collision,; see the inset in Fig. 8(b) and supplementary video 5. As seen in Figs. 8(b), 8(d), 8(f),and 8(h), formation of such clusters is indicated by an almost vanishing distance between the robots’ noses and the angle between the robots’ propulsion vectors tending to (the gray dashed lines in Fig. 8). In particular, during the first stage (lasting approximately s) of the experiment, the angle keeps its initial value , while the distance between robots’ noses decreases linearly. After that, the second stage of a collision takes place, hallmarked by an abrupt change in the angle from to accompanied by the distance reaching cm. Physically, this corresponds to the robots adjusting their propulsion vector directions by rotating around the point of contact until is reached. Finally, during the third stage, both and cm remain nearly unchanged (up to some noise-like fluctuations), highlighting that two robots now propel forward in a continuous contact. The described “heart” clusters essentially represent quarter-micelles and can potentially contribute to the formation of half-micelles (consisting of four robots) after a collision between two such clusters.
The described scenario of “heart” cluster formation is most clearly observed in Figs. 8(b) and 8(d) for the robots starting with which considerably differs from , while for the transition region is much less pronounced, as seen in Figs. 8(f) and 8(h). However, due to the increased technical difficulty in engineering the collisions between the robots for compared to , there are considerable deviations between the target value of and the actual impact angle obtained. In particular, Fig. 8(h) demonstrates close to in all three experiments. Moreover, in some cases collisions of the robots do not result in cluster formation [see the blue and orange solid lines in Fig. 8(h)].
In contrast, for -robots the “heart” clusters are unstable, as seen in Figs. 8(a) and 8(c). In particular, when and reach their values of and cm characteristic of the heart-shaped clusters, they fluctuate around their mean values for s. After that, starts growing until it reaches cm, while decreases to , as seen for all experiments in Fig. 8(a) and for the experiment shown with the green solid line in Fig. 8(c). The meaning of these characteristic values of and is uncovered in Fig. 8(g). It turns out that -robots support the formation of a different stable cluster type – the “tandem” cluster – in which one robot touches with its nose the round part of the other robot, as shown in the inset in Fig. 8(g) and supplementary video 5. Such clusters can also travel with the values of and retained, up to some fluctuations. Thus, the “heart” cluster in the case of -robots is inherently metastable and eventually transforms into the asymmetric “tandem” cluster described above. Note that for higher values of the impact angle , the formation of such “tandem” clusters is more pronounced for robots with abrasive lateral surfaces, as seen from the comparison of Fig. 8(e) and Fig. 8(g).
To summarize, the asymmetric “tandem” clusters observed in collisions of -robots do not represent quarter-micelles and are thus unlikely to contribute to micelle formation. In contrast, -robots form symmetric “heart” clusters that can potentially contribute to formation of micelles. Thus, the observed emergence of micellization appears to be related (at least partially) to two-particle effects, in particular to the formation of stable quarter-micelle motile clusters.
VI Discussion
In this work, we experimentally demonstrate the emergence of micellization in self-propelled particle ensembles implemented by robotic swarms, which is governed by an interplay of particles’ activity and their shape asymmetry, in contrast to micellization of surfactants in water solutions due to their inability to form hydrogen bonds Némethy and Scheraga (1962a, b, c). To prevent the dominance of boundary condensation Giomi et al. (2013); Deblais et al. (2018) which would inevitably mask bulk phenomena such as micelle formation, we consider two scenarios: systems of chiral robots moving on a flat surface along circular orbits with diameters much larger than the size of an individual robot, yet considerably smaller than the diameter of the boundary, and nonchiral robots moving linearly but placed in a parabolic satellite dish creating a soft localizing potential.
As we demonstrate, the presence or absence of micellization is defined by the location of the robot center of inertia, which supports the numerical predictions Kruglov and Borisov (2021). In particular, the robots can form micelles if their center of inertia is shifted towards the nose, while the micellization is suppressed for the center of inertia located close to the geometrical center. Such a behavior is clearly observed in both cases, for swarms of chiral robots on a flat surface and for swarms of nonchiral robots in a parabolic potential. The presence of such a localizing potential reduces the characteristic time of micelle formation and greatly enhances the mean lifetime of micelles. However, such micellelike clusters in the considered system can break apart due to the interaction with single robots or other clusters, and are generally metastable, in contrast to micelles formed by surfactants. As a result, a transition to the fully micellar phase in the considered system appears unrealistic. Additional studies of the stability of individual micelles as well as their interaction with the rest of the robotic swarm are discussed in the supplementary material Sup .
To underpin the microscopic physical mechanisms responsible for the formation of micelles on changing the location of the robot center of inertia, we performed two series of experiments addressing the motion characteristics of individual robots and the properties of two-robot collisions. Although the observed difference in the single-robot behavior cannot be directly linked to the emergence of micellization, our two-robot collision experiments clearly indicate the formation of two different types of stable two-robot clusters depending on the location of robot center of inertia. One of them represents a quarter of the micelle, thus being its constructive block. Therefore, the observed emergence of micellization is at least a two-particle effect.
The role of higher-order contributions and inertial effects Scholz et al. (2018); Leoni et al. (2020); Mandal et al. (2019) has not been considered in the present article, as arranging simultaneous three- or four-particle collisions in a laboratory setting and performing experiments to study collisions of individual robots with several-robot clusters appears to be a challenging task. It is a perspective direction for further study as it will allow one to uncover the role of many-particle microscopic effects in micelle formation and identify whether the two-particle processes considered in the present paper are the key factor. Moreover, various mechanisms might exist that can lead to a sufficient increase in the micelle stability that are mediated by the interaction of micelles and other clusters. For example, supplementary video 6 demonstrates an unusual herding of a complete micelle by a single robot roaming along the perimeter of the micelle that serves as a “shepherd” and continuously maintains the shape of the micelle, resulting in an anomalous lifetime of at least s (the cluster has formed at the timestamp s and remained stable throughout the experiment).
The second direction for further research is to consider swarms with nonzero net chirality in contrast to those studied in Sec. IV.2. For chiral particles, the emergence of edge states has been reported in physical settings involving swarms of various constituents ranging from self-rotating Yang et al. (2020) or self-propelled Barois et al. (2020) robots to bacteria with small net chirality Li et al. (2024) or even dancing pairs of people Du et al. (2024). In the context of robotic swarm micellization, the presence of such edge states could modify the rate of robot condensation at the boundary. Moreover, fully chiral micelles with a considerable angular momentum are expected to demonstrate stability properties and formation mechanisms different from those of nonchiral or slightly chiral micelles.
Development of an analytical theoretical description for the reported micellization is yet another direction for further research. The system considered in the present paper is inherently nonequilibrium as the robots constantly convert the energy stored in their batteries to kinetic energy. It is also not clear whether an effective temperature can be properly defined in such a system Loi et al. (2008); Takatori and Brady (2015); C.P. and Joy (2022); Boudet et al. (2022); Caprini et al. (2024). This renders the approach of considering micellization as a reversible chemical reaction at equilibrium Phillips (1955); Nagarajan (2014) (which is conventionally used to derive the free energy of micellization) inapplicable to our system. Moreover, in surfactant solutions, the micelle concentration increases as more surfactant is added to the mixture, in accordance with Le Chatelier’s principle. However, for the reported micellization, increasing the robot concentration in some cases instead leads to a decrease in the number of micellelike clusters observed for abrasive -robots, as is discussed in detail in the supplementary material Sup . Moreover, a further increase in the number of robots is expected to completely prevent micellization due to the jamming transition Barois et al. (2019).
Another prospective direction is the implementation of the reported micellization at smaller scales, for example, with millimeter-size particles propelled by Marangoni flows Shabaniverki et al. (2021); Kumar et al. (2022), and at the microscale, where new interaction mechanisms come into play, such as increased van der Waals forces. In terms of microscale implementations, Janus particles appear to be the most promising, as they allow one to change the activity of particles by an external illumination Villa and Pumera (2019); Palacci et al. (2013); Dai et al. (2016). Moreover, self-propelled Janus particles with various asymmetric shapes have been synthesized Dai et al. (2016); Shah et al. (2021); Liu et al. (2022), making the proposed design feasible.
Moreover, there is a wide range of possibilities in addressing complex external potentials Wang et al. (2021) and robot motion profiles Zhang et al. (2022) in order to develop novel control paradigms for robotic swarms Li et al. (2019), which can involve tools such as reinforcement learning Ben Zion et al. (2023). Finally, the reported micellization, once reproduced at the microscale, might find practical applications in the areas of microfluidics, microstructuring of materials, and waste removal. For example, micropumps based on self-rotating snowman-shaped particles have been demonstrated van den Beld et al. (2015). The use of self-propelled microparticles has also been proposed to synthesize self-organized random lasers Trivedi et al. (2022), microplastics removal Li et al. (2022); Won et al. (2022), and enhanced catalytic decomposition of medical masks Kim et al. (2023).
Acknowledgements.
We acknowledge fruitful discussions with Dr. Timofey Kruglov, who brought the idea of active - matter micellization to our attention; Dr. Alexander Borisov, Prof. Mikhail Lapine; and Prof. Anton Souslov. The work is supported by the Russian Science Foundation (project 24-79-10314).References
- Reynolds (1987) Craig W. Reynolds, “Flocks, herds and schools: A distributed behavioral model,” in Proceedings of the 14th annual conference on Computer graphics and interactive techniques - SIGGRAPH '87 (ACM Press, 1987).
- Vicsek et al. (1995) Tamás Vicsek, András Czirók, Eshel Ben-Jacob, Inon Cohen, and Ofer Shochet, “Novel type of phase transition in a system of self-driven particles,” Physical Review Letters 75, 1226–1229 (1995).
- Vicsek and Zafeiris (2012) Tamás Vicsek and Anna Zafeiris, “Collective motion,” Physics Reports 517, 71–140 (2012).
- Balasubramaniam et al. (2021) Lakshmi Balasubramaniam, Amin Doostmohammadi, Thuan Beng Saw, Gautham Hari Narayana Sankara Narayana, Romain Mueller, Tien Dang, Minnah Thomas, Shafali Gupta, Surabhi Sonam, Alpha S. Yap, Yusuke Toyama, René-Marc Mège, Julia M. Yeomans, and Benoît Ladoux, “Investigating the nature of active forces in tissues reveals how contractile cells can form extensile monolayers,” Nature Materials 20, 1156–1166 (2021).
- Pierce et al. (2018) C. J. Pierce, H. Wijesinghe, E. Mumper, B. H. Lower, S. K. Lower, and R. Sooryakumar, “Hydrodynamic interactions, hidden order, and emergent collective behavior in an active bacterial suspension,” Physical Review Letters 121, 188001 (2018).
- Be’er et al. (2020) Avraham Be’er, Bella Ilkanaiv, Renan Gross, Daniel B. Kearns, Sebastian Heidenreich, Markus Bär, and Gil Ariel, “A phase diagram for bacterial swarming,” Communications Physics 3, 66 (2020).
- Bricard et al. (2013) Antoine Bricard, Jean-Baptiste Caussin, Nicolas Desreumaux, Olivier Dauchot, and Denis Bartolo, “Emergence of macroscopic directed motion in populations of motile colloids,” Nature 503, 95–8 (2013).
- Giomi et al. (2013) L. Giomi, N. Hawley-Weld, and L. Mahadevan, “Swarming, swirling and stasis in sequestered bristle-bots,” Proceedings of the Royal Society A: Mathematical, Physical and Engineering Sciences 469, 20120637 (2013).
- Barois et al. (2019) Thomas Barois, Jean-François Boudet, Nicolas Lanchon, Juho Lintuvuori, and Hamid Kellay, “Characterization and control of a bottleneck-induced traffic-jam transition for self-propelled particles in a track,” Physical Review E 99, 052605 (2019).
- Patterson et al. (2020) German Patterson, D. Sornette, and Daniel Parisi, “Properties of balanced flows with bottlenecks: Common stylized facts in finance and vibration-driven vehicles,” Physical Review E 101, 042302 (2020).
- Li et al. (2021) Shengkai Li, Bahnsikha Dutta, Sarah Cannon, Joshua Daymude, Ram Avinery, Enes Aydin, André’a Richa, Daniel Goldman, and Dana Randall, “Programming active cohesive granular matter with mechanically induced phase changes,” Science Advances 7, eabe8494 (2021).
- Boudet et al. (2022) Jean François Boudet, Julie Jagielka, Thomas Guerin, Thomas Barois, Fabio Pistolesi, and Hamid Kellay, “Effective temperature and dissipation of a gas of active particles probed by the vibrations of a flexible membrane,” Physical Review Research 4, L042006 (2022).
- Cates and Tailleur (2015) Michael E. Cates and Julien Tailleur, “Motility-induced phase separation,” Annual Review of Condensed Matter Physics 6, 219–244 (2015).
- Palacci et al. (2013) Jeremie Palacci, Stefano Sacanna, Asher Preska Steinberg, David J. Pine, and Paul M. Chaikin, “Living crystals of light-activated colloidal surfers,” Science 339, 936–940 (2013).
- Ginot et al. (2018) F. Ginot, I. Theurkauff, F. Detcheverry, C. Ybert, and C. Cottin-Bizonne, “Aggregation-fragmentation and individual dynamics of active clusters,” Nature Communications 9, 696 (2018).
- Mousavi et al. (2019) S. Masoumeh Mousavi, Iryna Kasianiuk, Denis Kasyanyuk, Sabareesh K. P. Velu, Agnese Callegari, Luca Biancofiore, and Giovanni Volpe, “Clustering of janus particles in an optical potential driven by hydrodynamic fluxes,” Soft Matter 15, 5748–5759 (2019).
- Yang et al. (2020) Xiang Yang, Chenyang Ren, Kangjun Cheng, and H. P. Zhang, “Robust boundary flow in chiral active fluid,” Physical Review E 101, 022603 (2020).
- Shankar et al. (2022) Suraj Shankar, Anton Souslov, Mark J. Bowick, M. Cristina Marchetti, and Vincenzo Vitelli, “Topological active matter,” Nature Reviews Physics 4, 380–398 (2022).
- Souslov et al. (2017) Anton Souslov, Benjamin Zuiden, Denis Bartolo, and Vincenzo Vitelli, “Topological sound in active-liquid metamaterials,” Nature Physics 13, 1091–1094 (2017).
- Duclos et al. (2020) Guillaume Duclos, Raymond Adkins, Debarghya Banerjee, Matthew S. E. Peterson, Minu Varghese, Itamar Kolvin, Arvind Baskaran, Robert A. Pelcovits, Thomas R. Powers, Aparna Baskaran, Federico Toschi, Michael F. Hagan, Sebastian J. Streichan, Vincenzo Vitelli, Daniel A. Beller, and Zvonimir Dogic, “Topological structure and dynamics of three-dimensional active nematics,” Science 367, 1120–1124 (2020).
- Zhang et al. (2015) Qilu Zhang, Renfeng Dong, Xueyi Chang, Biye Ren, and Zhen Tong, “Spiropyran-decorated SiO2–pt janus micromotor: Preparation and light-induced dynamic self-assembly and disassembly,” ACS Applied Materials & Interfaces 7, 24585–24591 (2015).
- Kruglov and Borisov (2021) Timofei Kruglov and Alexander Borisov, “Micellisation and aggregation of self-propelled hard circulangles.” in Presentations and videos to 7th edition of the International Conference on Particle-based Methods (Particles 2021), Vol. CT07 (2021) p. 2.
- Scholz et al. (2021) Christian Scholz, Anton Ldov, Thorsten Pöschel, Michael Engel, and Hartmut Löwen, “Surfactants and rotelles in active chiral fluids,” Science Advances 7, eabf8998 (2021).
- Gütgemann et al. (2014) Ines Gütgemann, Hermann Heimpel, and Carl Thomas Nebe, “Significance of teardrop cells in peripheral blood smears,” LaboratoriumsMedizin 38, 000010151520140005 (2014).
- Dai et al. (2016) Baohu Dai, Jizhuang Wang, Ze Xiong, Xiaojun Zhan, Wei Dai, Chien-Cheng Li, Shien-Ping Feng, and Jinyao Tang, “Programmable artificial phototactic microswimmer,” Nature Nanotechnology 11, 1087–1092 (2016).
- Shah et al. (2021) Dylan Shah, Bilige Yang, Sam Kriegman, Michael Levin, Josh Bongard, and Rebecca Kramer-Bottiglio, “Shape changing robots: Bioinspiration, simulation, and physical realization,” Advanced Materials 33, e2002882 (2021).
- Liu et al. (2022) Meihuan Liu, Jiamiao Jiang, Haixin Tan, Bin Chen, Juanfeng Ou, Hong Wang, Jia Sun, Lu Liu, Fei Wang, Junbin Gao, Chang Liu, Fei Peng, Yun Liu, and Yingfeng Tu, “Light-driven Au–ZnO nanorod motors for enhanced photocatalytic degradation of tetracycline,” Nanoscale 14, 12804–12813 (2022).
- Dmitriev et al. (2023) Alexey A. Dmitriev, Alina D. Rozenblit, Vadim A. Porvatov, Mikhail K. Buzakov, Anastasia A. Molodtsova, Daria V. Sennikova, Vyacheslav A. Smirnov, Oleg I. Burmistrov, Timur I. Karimov, Ekaterina M. Puhtina, and Nikita A. Olekhno, “Swarmodroid 1.0: A modular bristle-bot platform for robotic active matter studies,” (2023), updated versions of the AMPy code and its documentation will be made available online (https://github.com/swarmtronics/ampy), arXiv:2305.13510 .
- Deblais et al. (2018) A. Deblais, T. Barois, T. Guerin, P.H. Delville, R. Vaudaine, J.S. Lintuvuori, J.F. Boudet, J.C. Baret, and H. Kellay, “Boundaries control collective dynamics of inertial self-propelled robots,” Physical Review Letters 120, 188002 (2018).
- Boudet et al. (2021) J. F. Boudet, J. Lintuvuori, C. Lacouture, T. Barois, A. Deblais, K. Xie, S. Cassagnere, B. Tregon, D. B. Brückner, J. C. Baret, and H. Kellay, “From collections of independent, mindless robots to flexible, mobile, and directional superstructures,” Science Robotics 6, abd0272 (2021).
- (31) See Supplemental Material at http://link.aps.org/supplemental/10.1103/4yzc-ppgy for (i) a description of the experimental setup; (ii) details on the choice of an abrasive material for robot side surfaces; (iii) detailed studies of a single-robot motion in a parabolic potential; (iv) experiments on a single-robot motion under a rotational excitation; (v) studies of two-robot collisions with different parameters; (vi) the classification of various several-robot clusters; (vii) studies of the effects of aggregation induced by a parabolic potential; and (viii) a comparison with micellization of surfactants. Additionally, Supplemental Video 1 demonstrates the condensation of robots at a vertical wall; Supplemental Video 2 demonstrates micellization at a flat surface with a vertical wall; Supplemental Video 3 demonstrates micellization in a parabolic potential; Supplemental Video 4 demonstrates the motion of a single robot in the parabolic potential; Supplemental Video 5 demonstrates the formation of stable two-robot clusters; Supplemental Video 6 demonstrates an example of a stabilized micelle; Supplemental Video 7 illustrates the stability of an isolated micelle like cluster; and Supplemental Video 8 shows the stability of artificially assembled micelle like cluster .
- Scholz et al. (2016) Christian Scholz, Sean D’Silva, and Thorsten Pöschel, “Ratcheting and tumbling motion of vibrots,” New Journal of Physics 18, 123001 (2016).
- Porvatov et al. (2021) Vadim A. Porvatov, Alina D. Rozenblit, Alexey A. Dmitriev, Oleg I. Burmistrov, Daria A. Petrova, Georgy Yu. Gritsenko, Ekaterina M. Puhtina, Egor I. Kretov, Dmitry S. Filonov, Anton Souslov, and Nikita A. Olekhno, “Optimizing self-rotating bristle-bots for active matter implementation with robotic swarms,” Journal of Physics: Conference Series 2086, 012202 (2021).
- Deseigne et al. (2010) Julien Deseigne, Olivier Dauchot, and Hugues Chaté, “Collective motion of vibrated polar disks,” Physical Review Letters 105, 098001 (2010).
- Kumar et al. (2014) Nitin Kumar, Harsh Soni, Sriram Ramaswamy, and A. Sood, “Flocking at a distance in active granular matter,” Nature Communications 5, 4688 (2014).
- Barois et al. (2020) Thomas Barois, Jean-François Boudet, Juho S. Lintuvuori, and Hamid Kellay, “Sorting and extraction of self-propelled chiral particles by polarized wall currents,” Physical Review Letters 125, 238003 (2020).
- Siebers et al. (2023) Frank Siebers, Ashreya Jayaram, Peter Blümler, and Thomas Speck, “Exploiting compositional disorder in collectives of light-driven circle walkers,” Science Advances 9, eadf5443 (2023).
- Ben Zion et al. (2023) Matan Yah Ben Zion, Jeremy Fersula, Nicolas Bredeche, and Olivier Dauchot, “Morphological computation and decentralized learning in a swarm of sterically interacting robots,” Science Robotics 8, eabo6140 (2023).
- Wang et al. (2021) Gao Wang, Trung V. Phan, Shengkai Li, Michael Wombacher, Junle Qu, Yan Peng, Guo Chen, Daniel I. Goldman, Simon A. Levin, Robert H. Austin, and Liyu Liu, “Emergent field-driven robot swarm states,” Physical Review Letters 126, 108002 (2021).
- Dauchot and Démery (2019) Olivier Dauchot and Vincent Démery, “Dynamics of a self-propelled particle in a harmonic trap,” Physical Review Letters 122, 068002 (2019).
- Hils et al. (2021) Christian Hils, Joachim Schmelz, Markus Drechsler, and Holger Schmalz, “Janus micelles by crystallization-driven self-assembly of an amphiphilic, double-crystalline triblock terpolymer,” Journal of the American Chemical Society 143, 15582–15586 (2021).
- Némethy and Scheraga (1962a) George Némethy and Harold A Scheraga, “Structure of water and hydrophobic bonding in proteins. I. A model for the thermodynamic properties of liquid water,” The Journal of Chemical Physics 36, 3382–3400 (1962a).
- Némethy and Scheraga (1962b) George Némethy and Harold A Scheraga, “Structure of water and hydrophobic bonding in proteins. II. Model for the thermodynamic properties of aqueous solutions of hydrocarbons,” The Journal of Chemical Physics 36, 3401–3417 (1962b).
- Némethy and Scheraga (1962c) George Némethy and Harold A Scheraga, “The structure of water and hydrophobic bonding in proteins. III. The thermodynamic properties of hydrophobic bonds in proteins,” The Journal of Physical Chemistry 66, 1773–1789 (1962c).
- Scholz et al. (2018) Christian Scholz, Soudeh Jahanshahi, Anton Ldov, and Hartmut Löwen, “Inertial delay of self-propelled particles,” Nature Communications 9, 5156 (2018).
- Leoni et al. (2020) Marco Leoni, Matteo Paoluzzi, Sarah Eldeen, Anthony Estrada, Lauren Nguyen, Maria Alexandrescu, Karin Sherb, and Wylie W. Ahmed, “Surfing and crawling macroscopic active particles under strong confinement: Inertial dynamics,” Physical Review Research 2, 043299 (2020).
- Mandal et al. (2019) Suvendu Mandal, Benno Liebchen, and Hartmut Löwen, “Motility-induced temperature difference in coexisting phases,” Physical Review Letters 123, 228001 (2019).
- Li et al. (2024) He Li, Hugues Chaté, Masaki Sano, Xia-qing Shi, and H. P. Zhang, “Robust edge flows in swarming bacterial colonies,” Physical Review X 14, 041006 (2024).
- Du et al. (2024) Matthew Du, Juan B Pérez-Sánchez, Jorge A Campos-Gonzalez-Angulo, Arghadip Koner, Federico Mellini, Sindhana Pannir-Sivajothi, Yong Rui Poh, Kai Schwennicke, Kunyang Sun, Stephan van den Wildenberg, et al., “Chiral edge waves in a dance-based human topological insulator,” Science Advances 10, eadh7810 (2024).
- Loi et al. (2008) Davide Loi, Stefano Mossa, and Leticia F Cugliandolo, “Effective temperature of active matter,” Physical Review E 77, 051111 (2008).
- Takatori and Brady (2015) Sho C. Takatori and John F. Brady, “Towards a thermodynamics of active matter,” Physical Review E 91, 032117 (2015).
- C.P. and Joy (2022) Sanjay C.P. and Ashwin Joy, “Effective temperature and einstein relation for particles in active matter flows,” Physical Review E 105, 065114 (2022).
- Caprini et al. (2024) Lorenzo Caprini, Davide Breoni, Anton Ldov, Christian Scholz, and Hartmut Löwen, “Dynamical clustering and wetting phenomena in inertial active matter,” Communications Physics 7, 343 (2024).
- Phillips (1955) J.N. Phillips, “The energetics of micelle formation,” Transactions of the Faraday Society 51, 561–569 (1955).
- Nagarajan (2014) Ramanathan Nagarajan, “One hundred years of micelles: evolution of the theory of micellization,” in Surfactant Science and Technology: Retrospects and Prospects, edited by Laurence S. Romsted (Taylor and Francis Boca Raton, FL, 2014) pp. 3–52.
- Shabaniverki et al. (2021) Soheila Shabaniverki, Antonio Alvarez-Valdivia, and Jaime J. Juárez, “3D printed self-propelled composite floaters,” Smart Materials and Structures 30, 075015 (2021).
- Kumar et al. (2022) Piyush Kumar, Yi Zhang, Stephen J. Ebbens, and Xiubo Zhao, “3d inkjet printed self-propelled motors for micro-stirring,” Journal of Colloid and Interface Science 623, 96–108 (2022).
- Villa and Pumera (2019) Katherine Villa and Martin Pumera, “Fuel-free light-driven micro/nanomachines: artificial active matter mimicking nature,” Chemical Society Reviews 48, 4966–4978 (2019).
- Zhang et al. (2022) Yifan Zhang, Renjie Zhu, Jianhao Wu, and Hongqiang Wang, “SimoBot: an underactuated miniature robot driven by a single motor,” IEEE/ASME Transactions on Mechatronics 27, 5748–5759 (2022).
- Li et al. (2019) Shuguang Li, Richa Batra, David Brown, Hyun-Dong Chang, Nikhil Ranganathan, Chuck Hoberman, Daniela Rus, and Hod Lipson, “Particle robotics based on statistical mechanics of loosely coupled components,” Nature 567, 361–365 (2019).
- van den Beld et al. (2015) Wesley TE van den Beld, Natalia L Cadena, Johan Bomer, Eddy L de Weerd, Leon Abelmann, Albert van den Berg, and Jan CT Eijkel, “Bidirectional microfluidic pumping using an array of magnetic Janus microspheres rotating around magnetic disks,” Lab on a Chip 15, 2872–2878 (2015).
- Trivedi et al. (2022) Manish Trivedi, Dhruv Saxena, Wai Ng, Riccardo Sapienza, and Giorgio Volpe, “Self-organized lasers from reconfigurable colloidal assemblies,” Nature Physics 18, 1–6 (2022).
- Li et al. (2022) Wanyuan Li, Changjin Wu, Ze Xiong, Chaowei Liang, Ziyi Li, Baiyao Liu, Qinyi Cao, Jizhuang Wang, Jinyao Tang, and Dan Li, “Self-driven magnetorobots for recyclable and scalable micro/nanoplastic removal from nonmarine waters,” Science Advances 8, eade1731 (2022).
- Won et al. (2022) Sukyoung Won, Hee Eun Lee, Young Shik Cho, Kijun Yang, Jeong Eun Park, Seung Jae Yang, and Jeong Jae Wie, “Multimodal collective swimming of magnetically articulated modular nanocomposite robots,” Nature Communications 13, 6750 (2022).
- Kim et al. (2023) Jeonghyo Kim, Carmen C Mayorga-Martinez, and Martin Pumera, “Magnetically boosted 1D photoactive microswarm for COVID-19 face mask disruption,” Nature Communications 14, 935 (2023).